### v1

@@ Abstract
Context:The kinematics and chemical composition of stellar populations of different ages provide crucial information on the evolution of the various components of a galaxy.
Aims:Our aim is to determine the kinematics of individual stars as a function of age in IC 1613, a star-forming, gas-rich and isolated dwarf galaxy of the Local Group.
Methods:We present results of a new spectroscopic survey of IC 1613 conducted with MUSE, an integral field spectrograph mounted on the Very Large Telescope. We extracted sources, from which we separated stellar objects for their subsequent spectral analysis. The quality of the dataset allowed us to obtain accurate classifications ( to better than K) and line-of-sight velocities (with average km s) for about 800 stars. Our sample includes not only red giant branch (RGB) and main sequence (MS) stars, but also a number of probable Be and C stars. We also obtained reliable metallicities ( dex) for about 300 RGB stars.
Results:The kinematic analysis revealed for the first time the presence of stellar rotation with high significance. We found general agreement with the velocity field of the neutral gas component, although on average the stars showed greater velocity dispersion and slower rotation, which can be attributed to asymmetric drift. Examining the kinematics of the different stellar components, the MS stars appear to follow closely that of the gas. The chemical analysis of the RGB stars revealed average properties comparable to other Local Group dwarf galaxies.
Conclusions:Our work represents a step forward in understanding the internal processes that govern the dynamical evolution of a low-mass galaxy.
From gas to stars: MUSEings on the internal evolution of IC 1613
@@ 1 Introduction
The ability to resolve stellar populations in nearby galaxies allows us to study the physical processes that govern their evolution. The dwarf galaxies of the Local Group (LG), in particular, provide an excellent laboratory due to their diverse characteristics, including size, gas content, star formation history (SFH), chemical abundances, and kinematics (e.g. Tolstoy et al. 2009; McConnachie 2012; Kirby et al. 2013; Gallart et al. 2015; Simon 2019; Putman et al. 2021; Battaglia & Nipoti 2022). Since these characteristics can be influenced by the proximity of a larger galaxy, such as the Milky Way and Andromeda, isolated LG dwarf galaxies provide valuable insights into the intrinsic mechanisms that regulate their evolution. In the gas-rich ones, the interplay between instability and stellar feedback processes can be reconstructed by combining the kinematics and spatial distribution of stars of different ages with the SFH and gas content of the galaxy (e.g. Leaman et al. 2017; Collins & Read 2022). Gathering this kind of information is usually an observational challenge, but the advent of integral field spectroscopy (IFS) has been transformative in this respect, due to the ability to inspect different galactic tracers with high observational efficiency (Roth et al. 2018, Sánchez et al. 2012; Cortese et al. 2014; Bundy et al. 2015; McLeod et al. 2020; Zoutendijk et al. 2020; Júlio et al. 2023; Vaz et al. 2023, but also).
@@
Among LG dwarf galaxies, IC 1613 is an ideal IFS target. On the one hand, it is a typical low-mass star-forming dwarf irregular galaxy (Skillman et al. 2014). On the other hand, stellar feedback processes have generated voids and bubbles in the interstellar medium whose impact on stellar and gas kinematics needs to be explored (Read et al. 2016). First discovered by Wolf (1906), IC 1613 was recognised soon afterwards as an extragalactic object by Baade (1935, as reviewed by Sandage 1971), who first determined its distance using Cepheid variable stars. Since then, several other measurements of its distance have been made using different indicators, including RR-Lyrae variable stars and the tip of the red giant branch (RGB). A compilation of literature distance values (including their own) was reported by Bernard et al. (2010), who obtained a statistical average value of , or kpc.
@@
IC 1613 therefore lies well beyond the Milky Way’s (MW) virial radius, as well as that of M31 (being at kpc distance from it, McConnachie 2012). The systemic proper motion of IC 1613 obtained from McConnachie et al. (2021) with Gaia DR2 data does not rule out the possibility that the galaxy was ever within 300 kpc of M31, for M31 masses of . However, the most recent estimate of the systemic proper motion of IC 1613 by Bennet et al. (2023), combining data from Gaia eDR3 and the Hubble Space Telescope (HST), makes such an association unlikely when assuming a mass of M31 of . Interestingly, Buck et al. (2019) assign to IC 1613 a high probability of being a backsplash galaxy (i.e. a currently isolated system that may have once passed close to a large host, but without becoming bound) of M31.
@@
Its low luminosity (an absolute magnitude of ) corresponds to a stellar mass of , similar to that of other LG dIrrs like NGC 6822 and WLM (McConnachie 2012). The SFH and stellar content of IC 1613 have been studied in detail over the years (see, e.g., Cole et al. 1999; Skillman et al. 2003; Bernard et al. 2007; Skillman et al. 2014). Based on deep HST Advanced Camera for Surveys (ACS) imaging, and after a review of archival HST data, Skillman et al. (2014) concluded that the SFH of IC 1613 has been close to constant on average throughout its life, with no evidence of an early dominant episode of star formation. We note that the HST/ACS field analysed by Skillman et al. (2014), roughly located at the half-light radius of the galaxy, was considered as representative of the global SFH of IC 1613. The comparison with archival HST observations located at different radii supported this assumption. The average metallicity was linearly increasing through time, with values ranging from dex to dex. Stars were formed at an average rate of .
@@
The structural properties of IC 1613 have been determined by studying the spatial distribution of stellar tracers in different evolutionary phases (see, e.g., Albert et al. 2000; Borissova et al. 2000; Bernard et al. 2007; Garcia et al. 2009; Sibbons et al. 2015; McQuinn et al. 2017; Pucha et al. 2019; Higgs et al. 2021). In general, it was found that the young stars are more centrally concentrated than the intermediate-age and old stars, a feature that is common in dwarfs. However, Pucha et al. (2019) using deep and wide Subaru/Hyper-SuprimeCam observations of IC 1613, showed that its young main sequence (MS) stars, along with the RGB and ancient horizontal branch (HB) stars, all extend to the outskirts of the galaxy, up to arcmin (i.e., effective radii ). In particular, the young stars are found well beyond the currently active star forming regions (within ), although with a much lower density compared to the intermediate and old age components. They also showed a steeper radial density profile in the inner regions than in the outer ones, in contrast to the RGB and HB stars. This seems to imply a different formation channel between the younger (e.g., from gas pushed outward through stellar feedback) and older stars (probably via gas accretion) in the galaxy’s outskirts. In this work, we assume the structural parameters from Higgs et al. (2021) who conducted an homogeneous analysis for the isolated LG dwarf galaxies.
@@
As other dIrrs, IC 1613 contains an extended component of neutral hydrogen (HI) gas, with a clumpy distribution on the inside, but showing regular contours at larger radii (see van den Bergh 2000, and references therein). In particular, the HI distribution is rich of shells and voids around the currently active star forming regions, where also the ionized gas is distributed (e.g., Lozinskaya et al. 2003; Silich et al. 2006; Moiseev & Lozinskaya 2012; Pokhrel et al. 2020). The HI kinematics shows a linearly increasing rotation curve with a km s (Lake & Skillman 1989; Oh et al. 2015; Read et al. 2016). The rotation value, however, may represent a lower limit due to the fact that this galaxy is probably seen face-on. This translates into a significant uncertainty in the determination of the dynamical mass being between and , depending on the assumed inclination angle (Read et al. 2016).
@@
The first spectroscopic measurements of stars in IC 1613 were obtained from some of the brightest ones, that is, from 9 early B-type young supergiants (Bresolin et al. 2007) and 3 evolved M-type supergiants (Tautvaišienė et al. 2007). Both studies obtained compatible metallicity values, dex and dex, respectively111Assuming dex and a solar oxygen abundance of dex (Asplund et al. 2009)., later confirmed by Berger et al. (2018) who studied 21 young BA-type supergiant stars. Interestingly, these authors found a bimodal metallicity distribution that appears to be correlated with their spatial location (i.e. they have lower metallicity in regions of high HI density). A comparable bi-modal distribution was also observed by Chun et al. (2022) for a sample of 14 red supergiants. Furthermore, there is evidence from the youngest stellar population and evolved red supergiants of IC 1613 that the present-day ratio could be sub-solar (i.e., dex, Tautvaišienė et al. 2007 and Garcia et al. 2014). The chemical abundance of the interstellar medium has been obtained spectroscopically from the central HII regions and is in good agreement with results from the supergiant stars (Bresolin et al. 2007, but see also Lee et al. 2003 and Tautvaišienė et al. 2007).
@@
Kirby et al. (2013) were the first to present a statistically significant spectroscopic analysis of a sample of 125 RGB stars in IC 1613, which led to an average metallicity of dex, in agreement with the fact that older stars are in general more metal-poor than the younger ones, and with expectations from the age-metallicity relation found by Skillman et al. (2014). On the same sample, Kirby et al. (2014) conducted a kinematic analysis which led to a systemic velocity of km s, in agreement with values from the HI component (Lake & Skillman 1989; Oh et al. 2015), and a velocity dispersion of km s. They did not find signs of rotation for the stellar component; this was later confirmed also by Wheeler et al. (2017) who re-analysed the Kirby et al. dataset. This is not surprising considering that the spatial distribution of this dataset is roughly along the optical minor axis of the galaxy, while in general it is expected that the stellar rotation follows that of the gas, which in this case rotates roughly along the optical major axis (Oh et al. 2015). The dynamical mass at the half-light radius reported by Kirby et al. (2014) was of .
@@
In this paper, we present a study of the stellar kinematic and chemical properties of IC 1613 from spectroscopic data taken with the MUSE integral field spectrograph on the Very Large Telescope (VLT). Taking advantage of the unique capabilities of this instrument, that combines high spatial resolution with a 1 arcmin field-of-view and a wide wavelength range at medium spectral resolution (Bacon et al. 2014), we performed a kinematic analysis as a function of stellar age and obtained metallicites for the largest spectroscopic stellar sample obtained to date for this galaxy. IC 1613 is located at high Galactic latitude ( deg), so its low values for both foreground (Schlafly & Finkbeiner 2011) and internal reddening (Georgiev et al. 1999) make it an ideal laboratory to compare the observed properties of its diverse stellar content.
@@
The article is structured as follows: in Sect. 2 we give details on the data acquisition and reduction processes we conducted. Section 3 is dedicated to spectral classification and velocity determination. In Sect. 4 we report details on the determination of likely member stars and their subsequent kinematic analysis. Section 5 shows results of the chemical analysis of the RGB stars. Finally, Sect. 6 is dedicated to the summary and perspectives of future work, while in the appendices we report details on the sanity checks conducted on the spectral-fitting codes used during our analysis. The parameters adopted for IC 1613 throughout the text are summarized in Table 1.
@@
2 Observations and Data Reduction
@@
Each field was observed with a total exposure time of 11080 s on source, split in s exposures, obtained by rotating the position angle of the spectrograph by 90 with respect to the previous exposure. No separate sky field was acquired since the stellar component of IC 1613 is resolved in the magnitude range of interest and the sky contribution is taken care of during the de-blending and spectral extraction phase of the data analysis (see Sect. 2.1). The data were acquired in dark time, clear sky, with a request for a seeing in V-band at zenith arcsec, fulfilled for all but one exposure of F1 (for which it deviated within 10% from the request). The observing log is reported in Table 7.
2.1 Data Reduction and Source Extraction
All data cubes were reduced using version 1.6 of the official MUSE pipeline (Weilbacher et al. 2012). The basic steps of the reduction cascade (bias subtraction, slice tracing, wavelength calibration, and basic processing) were performed using the default settings444This included correcting for the observer’s line of sight velocity with respect to the barycentre of the solar system.. This resulted in 24 pixel tables (one for each unit spectrograph) for each individual exposure. When combining the pixel tables for each exposure, flux calibration was also applied using standard stars observed on the same nights. Furthermore, the subtraction of sky emission lines was performed by determining their intensity directly from the scientific data, which contained sufficiently large patches of (almost) blank sky for this purpose. The sky emission lines were also used to quantify the quality of the wavelength calibration, which had an average accuracy of 0.03 Å, or 1.5 km s. The final step of the data reduction was the combination of all individual exposures obtained for a pointing. The end products of the data reduction were three data cubes, one for each field. Each cube has a dimension of approximately pixels.
@@ 2.1 Data Reduction and Source Extraction
PampelMUSE extracted 2293 initial spectra, examples of which are shown in Figs. 11, 12 and 13. For each extracted spectrum, PampelMUSE provided the associated flux uncertainty per pixel and an S/N value calculated around its central wavelength (hereafter S/N). However, as our sources cover a wide spectral range (see Sect. 3.1), we calculated two further S/N indicators: one around 5500 Å (S/N), and the other on the continuum around the CaT lines (S/N). This gave us a more accurate picture of the quality of our spectra depending on the spectral type of a given star.

### v2

@@ Abstract
Context. The kinematics and chemical composition of stellar populations of different ages provide crucial information on the evolution of the various components of a galaxy.
Aims. Our aim is to determine the kinematics of individual stars as a function of age in IC 1613, a star-forming, gas-rich and isolated dwarf galaxy of the Local Group (LG).
Methods. We present results of a new spectroscopic survey of IC 1613 conducted with MUSE, an integral field spectrograph mounted on the Very Large Telescope. We extracted sources, from which we separated stellar objects for their subsequent spectral analysis. The quality of the dataset allowed us to obtain accurate classifications ( to better than 500 K) and line-of-sight velocities (with average km s-1) for about 800 stars. Our sample includes not only red giant branch (RGB) and main sequence (MS) stars, but also a number of probable Be and C stars. We also obtained reliable metallicities ( dex) for about 300 RGB stars.
Results. The kinematic analysis of IC 1613 revealed for the first time the presence of stellar rotation with high significance. We found general agreement with the rotation velocity of the neutral gas component. Examining the kinematics of stars as a function of broad age ranges, we find that the velocity dispersion increases as a function of age, with the behaviour being very clear in the outermost pointings, while the rotation-to-velocity dispersion support decreases. On timescales of Gyr, the stellar kinematics still follow very closely that of the neutral gas, while the two components decouple on longer timescales. The chemical analysis of the RGB stars revealed average properties comparable to other Local Group dwarf galaxies. We also provide a new estimation of the inclination angle using only independent stellar tracers.
Conclusions. Our work provides the largest spectroscopic sample of an isolated LG dwarf galaxy. The results obtained seem to support the scenario in which the stars of a dwarf galaxy are born from a less turbulent gas over time.
@@ 1 Introduction
The ability to resolve stellar populations in nearby galaxies allows us to study the physical processes that govern their evolution. The dwarf galaxies of the Local Group (LG), in particular, provide an excellent laboratory due to their diverse characteristics, including size, gas content, star formation history (SFH), chemical abundances, and kinematics (e.g. Tolstoy et al. 2009; McConnachie 2012; Kirby et al. 2013; Gallart et al. 2015; Simon 2019; Putman et al. 2021; Battaglia & Nipoti 2022). Since these characteristics can be influenced by the proximity of a larger galaxy, such as the Milky Way and Andromeda, isolated LG dwarf galaxies provide valuable insights into the intrinsic mechanisms that regulate their evolution. In the gas-rich ones, the interplay between gravitational instabilities and stellar feedback processes can be reconstructed by combining the kinematics and spatial distribution of stars of different ages with the SFH and gas content of the galaxy (e.g. Leaman et al. 2017; Collins & Read 2022). Gathering this kind of information is usually an observational challenge, but the advent of integral field spectroscopy (IFS) has been transformative in this respect, due to the ability to inspect different galactic tracers with high observational efficiency (Roth et al. 2018, Sánchez et al. 2012; Cortese et al. 2014; Bundy et al. 2015; McLeod et al. 2020; Zoutendijk et al. 2020; Júlio et al. 2023; Vaz et al. 2023, but also).
@@
Among LG dwarf galaxies, IC 1613 is an ideal IFS target. On the one hand, it is a typical low-mass star-forming dwarf irregular galaxy (Skillman et al. 2014). On the other hand, stellar feedback processes have generated voids and bubbles in the interstellar medium whose impact on stellar and gas kinematics needs to be explored (Read et al. 2016). First discovered by Wolf (1906), IC 1613 was recognised afterwards as an extragalactic object by Baade (1935, as reviewed by Sandage 1971), who first determined its distance using Cepheid variable stars. Since then, several other measurements of its distance have been made using different indicators, including RR-Lyrae variable stars and the tip of the red giant branch (RGB). A compilation of literature distance values (including their own) was reported by Bernard et al. (2010), who obtained a statistical average value of , or kpc.
@@
IC 1613 therefore lies well beyond the Milky Way’s (MW) virial radius, as well as that of M31 (being at kpc distance from it, McConnachie 2012). The systemic proper motion of IC 1613 obtained from McConnachie et al. (2021) with Gaia DR2 data does not rule out the possibility that the galaxy was ever within 300 kpc of M31, for M31 virial masses of . However, the most recent estimate of the systemic proper motion of IC 1613 by Bennet et al. (2023), combining data from Gaia eDR3 and the Hubble Space Telescope (HST), makes such an association unlikely when assuming an M31 virial mass of . Interestingly, Buck et al. (2019) assign to IC 1613 a high probability of being a backsplash galaxy (i.e. a currently isolated system that may have once passed close to a large host, but without becoming bound) of M31.
@@
Its low luminosity (an absolute magnitude of mag) corresponds to a total stellar mass of , similar to that of other LG dIrrs like NGC 6822 and WLM (McConnachie 2012). The SFH and stellar content of IC 1613 have been studied in detail over the years (see e.g. Cole et al. 1999; Skillman et al. 2003; Bernard et al. 2007; Skillman et al. 2014). Based on deep HST Advanced Camera for Surveys (ACS) imaging, and after a review of archival HST data, Skillman et al. (2014) concluded that the SFH of IC 1613 has been close to constant on average throughout its life, with no evidence of an early dominant episode of star formation. We note that the HST/ACS field analysed by Skillman et al. (2014), roughly located at the half-light radius of the galaxy, was considered as representative of the global SFH of IC 1613. The comparison with archival HST observations located at different radii supported this assumption. The average metallicity was linearly increasing through time, with values ranging from dex to dex. Stars were formed at an average rate of .
@@
The structural properties of IC 1613 have been determined by studying the spatial distribution of stellar tracers in different evolutionary phases (see, e.g., Albert et al. 2000; Borissova et al. 2000; Bernard et al. 2007; Garcia et al. 2009; Sibbons et al. 2015; McQuinn et al. 2017; Pucha et al. 2019; Higgs et al. 2021). In general, it was found that the young stars are more centrally concentrated than the intermediate-age and old stars, a feature that is common in dwarfs. However, Pucha et al. (2019) using deep and wide Subaru/Hyper-SuprimeCam observations of IC 1613, showed that its young main sequence (MS) stars, along with the RGB and ancient horizontal branch (HB) stars, all extend to the outskirts of the galaxy, up to arcmin (i.e., effective radii ). In particular, the young stars are found well beyond the currently active star forming regions (within ), although with a much lower density compared to the intermediate and old age components. They also showed a steeper radial density profile in the inner regions than in the outer ones, in contrast to the RGB and HB stars. This seems to imply a different formation channel between the younger (e.g., from gas pushed outward through stellar feedback) and older stars (probably via gas accretion) in the galaxy’s outskirts. In this work, we assume the structural parameters from Higgs et al. (2021) who conducted an homogeneous analysis of the isolated LG dwarf galaxies.
@@
As other dIrrs, IC 1613 contains an extended component of neutral hydrogen (HI) gas, with a clumpy distribution on the inside, but showing regular contours at larger radii (see van den Bergh 2000, and references therein). In particular, the HI distribution is rich of shells and voids around the currently active star forming regions, where also the ionized gas is distributed (e.g., Lozinskaya et al. 2003; Silich et al. 2006; Moiseev & Lozinskaya 2012; Pokhrel et al. 2020; Yarovova et al. 2024). The HI kinematics shows a linearly increasing rotation curve with a km s-1 (Lake & Skillman 1989; Oh et al. 2015; Read et al. 2016). The rotation value, however, may represent a lower limit due to the fact that this galaxy is probably seen close to face-on. This translates into a significant uncertainty in the determination of the dynamical mass being between and , depending on the assumed inclination angle (Read et al. 2016).
@@
The first spectroscopic measurements of stars in IC 1613 were obtained from some of the brightest ones, that is, from 9 early B-type young supergiants (Bresolin et al. 2007) and 3 evolved M-type supergiants (Tautvaišienė et al. 2007). Both studies obtained compatible metallicity values, dex and dex, respectively111Assuming dex and a solar oxygen abundance of dex (Asplund et al. 2009)., later confirmed by Berger et al. (2018) who studied 21 young BA-type supergiant stars. Interestingly, these authors found a bimodal metallicity distribution that appears to be correlated with their spatial location (i.e. lower metallicity stars are found in high density HI regions). A comparable bi-modal distribution was also observed by Chun et al. (2022) for a sample of 14 red supergiants. Furthermore, there is evidence from the youngest stellar population and evolved red supergiants of IC 1613 that the present-day ratio could be sub-solar (i.e., dex, Tautvaišienė et al. 2007 and Garcia et al. 2014). The chemical abundance of the interstellar medium has been obtained spectroscopically from the central HII regions and is in good agreement with results from the supergiant stars (Bresolin et al. 2007, but see also Lee et al. 2003 and Tautvaišienė et al. 2007).
@@
Kirby et al. (2013) were the first to present a statistically significant spectroscopic analysis of a sample of 125 RGB stars in IC 1613, which led to an average metallicity of dex, in agreement with the fact that older stars are in general more metal-poor than the younger ones, and with expectations from the age-metallicity relation found by Skillman et al. (2014). On the same sample, Kirby et al. (2014) conducted a kinematic analysis which led to a systemic velocity of km s-1, in agreement with values from the HI component (Lake & Skillman 1989; Oh et al. 2015), and a velocity dispersion of km s-1. They did not find signs of rotation for the stellar component; this was later confirmed also by Wheeler et al. (2017) who re-analysed the Kirby et al. dataset. This is not surprising considering that the spatial distribution of this dataset is roughly perpendicular to the HI major kinematic axis (Lake & Skillman 1989; Oh et al. 2015), assuming that the stellar rotation follows that of neutral gas. The dynamical mass at the half-light radius reported by Kirby et al. (2014) was .
@@
In this paper, we present a study of the stellar kinematic and chemical properties of IC 1613 from spectroscopic data taken with the MUSE integral field spectrograph on the Very Large Telescope (VLT). We took advantage of the unique capabilities of this instrument, that combines high spatial resolution with a 1 arcmin2 field-of-view and a wide wavelength range at medium spectral resolution (Bacon et al. 2014). We then performed a kinematic analysis as a function of stellar age and obtained metallicities for the largest spectroscopic stellar sample obtained to date for this galaxy. IC 1613 is located at high Galactic latitude ( deg), so its low values for both foreground (Schlafly & Finkbeiner 2011) and internal reddening (Georgiev et al. 1999) make it an ideal laboratory to compare the observed properties of its diverse stellar content.
@@
The article is structured as follows: in Sect. 2 we give details on the data acquisition and reduction processes we conducted. Section 3 is dedicated to spectral classification and velocity determination. In Sect. 4 we report details on the determination of likely member stars and their subsequent kinematic analysis, while in Sect. 5 we show results of the chemical analysis of the RGB stars in our sample. Section 6 is dedicated to the discussion of the age-kinematic trends of the different tracers inspected in this work. In this section, we also discuss the dynamical mass estimation and the role played by the galaxy’s inclination angle. Finally, Sect. 7 is dedicated to the summary and perspectives of future work, while in the appendices we report details on the sanity checks conducted during our analysis. The parameters adopted for IC 1613 throughout the text are summarised in Table 1.
@@
2 MUSE data of IC 1613
2.1 Observations and data reduction
@@
Each field was observed with a total exposure time of 11080 s on source, split in s exposures, obtained by rotating the position angle of the spectrograph by 90∘ with respect to the previous exposure. No separate sky field was acquired since the stellar component of IC 1613 is resolved in the magnitude range of interest and the sky contribution is taken care of during the de-blending and spectral extraction phase of the data analysis (see Sect. 2.2). The data were acquired in dark time, clear sky, with a request for a seeing in V-band at zenith arcsec, fulfilled for all but one exposure of F1 (for which it deviated within 10% from the request). The observing log is reported in Table 7.
All data cubes were reduced using version 1.6 of the official MUSE pipeline (Weilbacher et al. 2012). The basic steps of the reduction cascade (bias subtraction, slice tracing, wavelength calibration, and basic processing) were performed using the default settings444This included correcting for the observer’s line of sight velocity with respect to the barycentre of the solar system.. This resulted in 24 pixel tables (one for each unit spectrograph) for each individual exposure. When combining the pixel tables for each exposure, flux calibration was also applied using standard stars observed on the same nights. Furthermore, the subtraction of sky emission lines was performed by determining their intensity directly from the scientific data, which contained sufficiently large patches of (almost) blank sky for this purpose. The sky emission lines were also used to quantify the quality of the wavelength calibration, which had an average accuracy of 0.03 Å, or 1.5 km s-1. The final step of the data reduction was the combination of all individual exposures obtained for a pointing. The end products of the data reduction were three data cubes, one for each field. Each cube has a dimension of approximately pixels. In Fig. 2, we show the broad- and narrow-band images obtained from the reduced data cubes.
2.2 Source extraction and photometric cleaning
@@
PampelMUSE extracted 2293 initial spectra, examples of which are shown in Figs. 12, 13 and 14. For each extracted spectrum, PampelMUSE provided the associated flux uncertainty per pixel and an S/N value calculated around its central wavelength (hereafter S/NC). However, as our sources cover a wide spectral range (see Sect. 3.1.1), we calculated two further S/N indicators: one around 5500 Å (S/N550), and the other on the continuum around the CaT lines (S/NCaT). This gave us a more accurate picture of the quality of our spectra depending on the spectral type of a given star.